OpenStudy (anonymous):

choose the point-slope form of the equation below that represents the line that passes through the points (-6, 4) and (2, 0). y - 4 + - 1/2 (x + 6) y - 4 + 2 (x - 4) y + 6 + -1/2 (x - 4) y + 6 = 2 (x - 4)

OpenStudy (anonymous):

\[y-4=\frac{ 0-4 }{ 2+6 }*(x+6)\] y-4=-1/2(x+6)

OpenStudy (anonymous):

option A there is some error in your question at the place of + it will be= miss toxic ;)

OpenStudy (anonymous):

oops, lol thanks :)

OpenStudy (anonymous):

keep rocking ;)

OpenStudy (anonymous):

rock solid

OpenStudy (anonymous):

yayyyy lol